### v1

@@ 7 Internal structure
Using the masses and transit depths reported in Table 2, we ran the neural network based internal structure modelling framework plaNETic212121https://github.com/joannegger/plaNETic (Egger et al., 2024) to infer the internal structure of TOI-396 b and d. plaNETic uses a full grid accept-reject sampling method in combination with a deep neural network (DNN) that was trained on the forward model of BICEPS (Haldemann et al., 2024) to infer the internal structure of observed planets. Each planet is modelled as a three-layered structure: an inner iron-dominated core, a silicate mantle and a fully mixed envelope made up of water and H/He. In the case of multi-planet systems, all planets are modelled simultaneously.
@@
Figures 10 and 11 show the resulting posteriors of the most important interior structure parameters for TOI-396 b and d, respectively, in comparison with the chosen priors (dotted lines). Tables 9 and 10 in the Appendix summarise the median and one sigma error intervals for the full set of internal structure parameters. For both planets, the posterior distributions for the core and mantle mass fractions largely agree with the chosen priors for each of the six models that we ran. Indeed, the only planetary structure parameters for which the observational data contributed to their characterisation were the envelope mass fractions. If we assume that the planets formed outside the iceline, we find envelope mass fractions of % (A1), % (A2) and % (A3) for planet b and % (A1), % (A2) and % (A3) for planet d, with water mass fractions in the envelope of almost 100%. Conversely, if the planets were to have formed inside the iceline, we infer envelope mass fractions of the order of for planet b and for planet d, almost entirely made up of H/He.
@@ 8 JWST characterisation prospects
Given the bright host star and combination of planetary masses, radii, and equilibrium temperatures, the three planets have favourable metrics for atmospheric characterisation in both transmission and emission among sub-Neptunes (Kempton et al., 2018, see Fig. 12). This makes the TOI-396 system a highly valuable laboratory to study the formation and evolution of planetary systems. Thus, we explored the prospects for characterisation with JWST. We focused these simulations on emission observations, but note that either transmission or emission have their own advantages and disadvantages in terms of achievable science goals and challenges.
@@ 9 Conclusions
TOI-396 is an F6 V bright naked-eye star orbited by three planets of almost equal size, where the two inner ones are close to-, but out of- a 5:3 MMR. A photometric analysis of the system was already performed by Vanderburg et al. (2019), but by benefiting of two further TESS sectors we improved the precision on the planet radii by a factor 1.4, obtaining , , and .
@@
We determined the planet masses by extracting the RV time series from HARPS CCFs using an SN fit followed by a joint LC & RV MCMC analysis, where the RV detrending uses the breakpoint method. We get a firm detection of the RV signals of planets b and d, deriving and , while we can provide only a 3 upper limit for the mass of TOI-396 c of . This yields the following mean planet densities: , , and g cm-3 implying a quite unusual system architecture (Mishra et al., 2023) with the mid planet being the least dense and the outermost planet being the densest.
@@
The reason for the RV non-detection of any Keplerian signal at 6 d is likely to be ascribed to the vicinity of to the stellar rotation period. As a matter of fact, from the GLS periodograms of both the RV-related activity indices and the TESS raw LCs and from -based empirical relations we consistently infer d. After injecting synthetic Keplerian signals at and different semi-amplitudes () into the RV time series, we empirically find that the RV semi-amplitudes outputted by the MCMC analyses () are systematically lower than the input ones by almost 3 and they are statistically non-significant as far as . In addition, we find that when considering a planet with 3 (i.e. 2 g cm-3), which might correspond to the properties of TOI-396 c. On a more general perspective, these simulations confirm that stellar activity destructively interferes with Keplerian signals having (e.g. Vanderburg et al., 2016) and, furthermore, they indicate that – even in case of firm detection – are significantly underestimated.
@@
Longer-baseline RV observations may help disentangle coherent signals originated by Keplerian motions from non-coherent signals due to stellar activity, even if degeneracy issues still hold when the planet orbital period is close to the stellar rotation period (Kossakowski et al., 2022). Alternatively, possible constraint on may come from TTVs, as planets b and c are close to a MMR of the second order. Indeed, the TTV amplitudes of the two planets show a characteristic anti-correlation pattern as expected, however the phase coverage given by the available observations is too poor to perform a conclusive TTV dynamical analysis based on the observed transit timings of the planets. We also attempted to fit the TTV and RV simultaneously while integrating the orbits of the system. We found that the masses and densities of planets b and d are consistent with the results from the joint LC & RV analysis. TOI-396 c shows a dynamical mass of , greater than that inferred from the joint LC & RV analysis, but consistent (); the density is consistent at the level. However, we emphasise that, although formally precise, the estimate might not be accurate, as the full coverage of the TTV phase is needed to reliably compute TTV-based masses. Therefore, to fully confirm the system architecture we still miss a reliable estimate of the mass of TOI-396 c.
@@
We also checked the evolution of the system over 10 000 years and the critical resonance angles showed that planets b and c are close to, but not in, 5:3 MMR. We further performed forward N-body simulations over a temporal baseline of 5.2 years to track the transit epochs and evaluate the expected TTV amplitudes during time. It turns out that TOI-396 b and TOI-396 c may exhibit TTVs with a super-period of about 5 years and semi-amplitudes of 2 and 5 hours, respectively. This translates into a temporal drift of the transit timings that can raise up to 5 and 10 hours with respect to the linear ephemerides computed from TESS data.
@@
Characterising the nature of the planetary atmosphere is also key to correctly assess the planetary bulk densities (in particular for planet c). The potentially high TTVs inferred from our simulations should be duly taken into account when scheduling future observations of the target. This holds not only for JWST, but also for CHEOPS (Benz et al., 2021) that looks particularly suitable for collecting exquisite photometric data to enable the full characterisation of the system.
@@ Acknowledgements.
- Sousa (2014) Sousa, S. G. 2014, [arXiv:1407.5817] [arXiv:1407.5817]
- Sousa et al. (2021) Sousa, S. G., Adibekyan, V., Delgado-Mena, E., et al. 2021, arXiv e-prints, arXiv:2109.04781
@@ Appendix A Supplementary material
1]

### v2

@@ 7 Internal structure
Using the masses and transit depths reported in Table 2, we ran the neural network based internal structure modelling framework plaNETic232323https://github.com/joannegger/plaNETic (Egger et al., 2024) to infer the internal structure of TOI-396 b and d. plaNETic uses a full grid accept-reject sampling method in combination with a deep neural network (DNN) that was trained on the forward model of BICEPS (Haldemann et al., 2024) to infer the internal structure of observed planets. Each planet is modelled as a three-layered structure: an inner iron-dominated core, a silicate mantle and a fully mixed envelope made up of water and H/He. In the case of multi-planet systems, all planets are modelled simultaneously.
@@
Figures 10 and 11 show the resulting posteriors of the most important interior structure parameters for TOI-396 b and d, respectively, in comparison with the chosen priors (dotted lines). Tables 4 and 5 in the Appendix summarise the median and one sigma error intervals for the full set of internal structure parameters. For both planets, the posterior distributions for the core and mantle mass fractions largely agree with the chosen priors for each of the six models that we ran. Indeed, the only planetary structure parameters for which the observational data contributed to their characterisation were the envelope mass fractions. If we assume that the planets formed outside the iceline, we find envelope mass fractions of % (A1), % (A2) and % (A3) for planet b and % (A1), % (A2) and % (A3) for planet d, with water mass fractions in the envelope of almost 100%. Conversely, if the planets were to have formed inside the iceline, we infer envelope mass fractions of the order of for planet b and for planet d, almost entirely made up of H/He.
@@ 8 JWST characterisation prospects
Given the bright host star and combination of planetary masses, radii, and equilibrium temperatures, the three planets have favourable metrics for atmospheric characterisation in both transmission and emission among sub-Neptunes (Kempton et al., 2018, see Fig. 12). This makes the TOI-396 system a highly valuable laboratory to study the formation and evolution of planetary systems. Thus, we explored the prospects for characterisation with JWST. We focused these simulations on emission observations, but we note that transmission and emission have their own advantages and disadvantages in terms of achievable science goals and challenges.
@@ 9 Conclusions
The object TOI-396 is an F6 V bright naked-eye star orbited by three planets of almost equal size, and the two inner planets are close to but out of a 5:3 MMR. A photometric analysis of the system was already performed by Vanderburg et al. (2019), but by benefiting from two additional TESS sectors, we improved the precision on the planet radii by a factor of 1.4, obtaining , , and .
@@
We determined the masses of the planets by extracting the RV time series from HARPS CCFs using an SN fit followed by a joint LC and RV MCMC analysis, where the RV de-trending uses the breakpoint method. We obtained a firm detection of the RV signals of planets b and d, deriving and , but we can provide only a 3 upper limit for the mass of TOI-396 c of . This yields the following mean planet densities: , , and g cm-3, implying a quite unusual system architecture (Mishra et al., 2023) where the mid planet is the least dense and the outermost planet is the densest.
@@
The reason for the RV non-detection of any Keplerian signal at 6 d is likely to be ascribed to the vicinity of to the stellar rotation period. As a matter of fact, from the GLS periodograms of both the RV-related activity indices and the TESS raw LCs and from -based empirical relations, we consistently inferred d. After injecting synthetic Keplerian signals at and different semi-amplitudes () into the RV time series, we empirically find that the RV semi-amplitudes output by the MCMC analyses () are systematically lower than the input ones by almost 3, and they are statistically non-significant as far as . In addition, we find that when considering a planet with 3 (i.e. 2 g cm-3), which might correspond to the properties of TOI-396 c. On a more general perspective, these simulations confirm that stellar activity destructively interferes with Keplerian signals having (e.g. Vanderburg et al., 2016), and furthermore, they indicate that – even in the case of firm detection – values of are significantly underestimated.
@@
Longer-baseline RV observations may help disentangle coherent signals originated by Keplerian motions from non-coherent signals due to stellar activity, even if degeneracy issues still hold when the planet orbital period is close to the stellar rotation period (Kossakowski et al., 2022). Alternatively, a possible constraint on may come from TTVs, as planets b and c are close to an MMR of the second order. Indeed, the TTV amplitudes of the two planets show a characteristic anti-correlation pattern, as expected; however, the phase coverage given by the available observations is too poor to perform a conclusive TTV dynamical analysis based on the observed transit timings of the planets. We also attempted to fit the TTV and RV simultaneously while integrating the orbits of the system. We found that the masses and densities of planets b and d are consistent with the results from the joint LC and RV analysis. TOI-396 c shows a dynamical mass of , which is greater than that inferred from the joint LC and RV analysis, but it is consistent (); the density is consistent at the level. However, we emphasise that, although formally precise, the estimate might not be accurate, as the full coverage of the TTV phase is needed to reliably compute TTV-based masses. Therefore, to fully confirm the system architecture, a reliable estimate of the mass of TOI-396 c is still missing.
@@
We also checked the evolution of the system over 10 000 years, and the critical resonance angles showed that planets b and c are close to but not in a 5:3 MMR. We further performed forward N-body simulations over a temporal baseline of 5.2 years in order to track the transit epochs and evaluate the expected TTV amplitudes during time. It turns out that TOI-396 b and TOI-396 c may exhibit TTVs with a super-period of about 5 years and semi-amplitudes of 2 and 5 hours, respectively. This translates into a temporal drift of the transit timings that can rise up to 5 and 10 hours with respect to the linear ephemerides computed from TESS data.
@@
Characterising the nature of the planetary atmosphere is also key to correctly assessing the planetary bulk densities (in particular for planet c). The potentially high TTVs inferred from our simulations should be duly taken into account when scheduling future observations of the target. This holds not only for JWST, but also for CHEOPS (Benz et al., 2021), which appears especially suitable for collecting exquisite photometric data to enable the full characterisation of the system.
@@ Acknowledgements.
- Sousa (2014) Sousa, S. G. 2014, in Determination of Atmospheric Parameters of B-, A-, F- and G-Type Stars., ed. E. Niemczura, B. Smalley, & W. Pych (Springer International Publishing (Cham)), 297–310
- Sousa et al. (2021) Sousa, S. G., Adibekyan, V., Delgado-Mena, E., et al. 2021, A&A, 656, A53
@@ Appendix A Supplementary material
![[Uncaptioned image]](x8.png)
![[Uncaptioned image]](x9.png)
![[Uncaptioned image]](x10.png)
![[Uncaptioned image]](x11.png)
![[Uncaptioned image]](x12.png)
![[Uncaptioned image]](x13.png)
![[Uncaptioned image]](x14.png)